OpenStudy (anonymous):

how do you graph 5x + 2y = 10

OpenStudy (lgbasallote):

are you familiar with the intercept method?

OpenStudy (anonymous):

ya thats actually what im learning in class, but i missed a few days so it's really confusing.

OpenStudy (lgbasallote):

it's very basic and can be taught in a few seconds. to find the x-intercept: let y = 0 then solve for x for example x + 2y = 4 if i let y = 0 x + 2(0) = 4 x + 0 = 4 x = 4 so the x-intercept is (4,0) to find the y-intercept: let x = 0 then solve for y x + 2y = 4 if i let x = 0 0 + 2y = 4 2y = 4 y = 4/2 y = 2 so the y-intercept is (0,2) did you follow that?

OpenStudy (anonymous):

slightly, but that way of explaining is almost the exact same words as my textbook.

OpenStudy (lgbasallote):

well...they are the official wording...anyway...the best way for you to learn is for you to try it yourself 5x + 2y = 10 solve for x-intercept first

OpenStudy (anonymous):

5x + 0 = 10 5x = 10 x = 2 0 + 2y = 10 2y = 10 y = 10/2 y = 5 x = 2 5 times x = 10 2 times y = 10 10 + 10 doesn't equal 10 where did i go wrong?

OpenStudy (lgbasallote):

when you tried checking if you were right

OpenStudy (lgbasallote):

y = 5 and x = 2 are right HOWEVER y = 5 WHEN x= 0 (remember you set this to 0 earlier) x = 2 WHEN y = 0 (you set y to 0)

OpenStudy (lgbasallote):

therefore, the way to check is definitely not by substitution

OpenStudy (lgbasallote):

anyway... you have x= 2 and y = 5 therefore, the intercepts are (2,0) and (0,5) agree?

OpenStudy (anonymous):

ohhhhhh yes, thanks

OpenStudy (lgbasallote):

so now, all that's left is to plot (2,0) and (0,5) then connect the dots